A bicycle travels up a hill at an average speed of 4m/s and down the same hill through the same distance at an average speed of 8m/s. Express the average speed for the entire journey in terms oftandd.

Grade:12

2 Answers

RAVI KANT KUMAR
10 years ago
avg speed=total dist/total time let s = dist covered total dist=2s total time=s/4 + s/8 = 3s/8 avg speed=16/3 m/s =5.33 m/s
raj
10 years ago
You have to speculate given this convoluted question. Let us assume that the time taken to travel up the hill istseconds. Then, the time taken to travel down the hill is 1/2 oftseconds as the speed now is twice than that of the former. The total distance traversed, we assume, is 2d metres. Hence, the average speed is:2d/(1/2 + 1)t.There you go. You might have realized that this question isn`t testing you on physics; rather, it is testing you on math.